# Spectral Small-Incremental-Entangling: Breaking Quasi-Polynomial Complexity Barriers in Long-Range Interacting Systems

### v1

@@ III.3 Spectral-entangling strength
A central problem in this work is efficiently characterizing the capacity of an operator to generate entanglement spectra across subsystems. However, within existing methodologies, there has not been an established quantity that efficiently and directly quantifies the ability of a general operator to generate the entanglement spectrum. To address this gap, we introduce a new quantity, which we refer to as the Spectral Entangling (SE) Strength.
@@
Definition 1 (Spectral Entangling (SE) Strength).
@@ IV.1 Decay of the Schmidt Coefficients
$\langle A_{j^{\prime}}|A_{j}\rangle\neq 0,\quad\langle B_{j^{\prime}}|B_{j}\rangle\neq 0\quad\textrm{for}\quad j\neq j^{\prime}$ (21)
@@ IV.1.1 Proof of Lemma 1
Proof.
Using Eq. (29) for each $\lambda_{s}(\phi)$, and applying the definition of the SE strength, we deduce the inequality of
@@
Taking the square root on both sides yields the desired bound. ∎
@@ V.1 Entanglement rate for Rényi entanglement ($\alpha\geq 1/2$)
$E_{\alpha}(\psi_{t})=\frac{1}{1-\alpha}\log\left[{\rm tr}_{A}\left(\rho_{A}(t)^{\alpha}\right)\right],\quad\rho_{A}(t)={\rm tr}_{B}\left(|\psi_{t}\rangle\langle\psi_{t}|\right),$ (49)
we obtain the upper bound of the entanglement rate as
@@
The statement remains valid even when the system is extended by attaching an arbitrary ancilla to $A$ and $B$.
@@
where we used the inequality (18) to obtain $\bar{{\mathcal{J}}}(V_{AB})\leq\sum_{j}|J_{j,AB}|$.
@@ V.2 Optimality of the entanglement generation
Before showing the proof, we emphasize the optimality of the entanglement rate bound for the Rényi entanglement. The main points are listed below:
@@
As additional remarks on optimality, we also note the following:
@@ V.3 Proof of Theorem 1
$\frac{d}{dt}E_{\alpha}(\psi_{t})=\frac{1}{1-\alpha}\frac{d}{dt}\log\left[{\rm tr}_{A}\left(\rho_{t,A}^{\alpha}\right)\right]=\frac{\alpha}{(1-\alpha){\rm tr}\left(\rho_{t,A}^{\alpha}\right)}{\rm tr}_{AB}\left(\rho_{t,A}^{\alpha-1}[-iH,\rho_{t}]\right)$ (59)
where $\rho_{t}=|\psi_{t}\rangle\langle\psi_{t}|$. Hence, we obtain
@@
In the following, we generally consider the upper bound of $\left|{\rm tr}_{AB}\left(\rho_{A}^{\alpha-1}[H,\rho]\right)\right|$ for an arbitrary quantum state $\rho=|\phi\rangle\langle\phi|$.
@@
$\tilde{c}_{\alpha}:=\left|\left(2\alpha-1\right)^{1/(2-2\alpha)}-\left(2\alpha-1\right)^{(2\alpha-1)/(2-2\alpha)}\right|.$ (64)
Note that $\left|x-x^{2\alpha-1}\right|$ attains its maximum when $x=(2\alpha-1)^{1/(2-2\alpha)}$.
@@
From Definition (1) for the SE strength, $V_{AB}|\phi_{s,A},\phi_{s,B}\rangle$ can be rewritten as
@@ V.4 Tightness at $\alpha=1/2$
$\left|\frac{E_{1/2}(\psi_{t})-E_{1/2}(\psi_{0})}{t}\right|\geq 2\bar{{\mathcal{J}}}(V_{A_{0}B_{0}})-\frac{2[t\bar{{\mathcal{J}}}(V_{A_{0}B_{0}})]^{2}}{n}$ (76)
@@
In the inequality (76), only the first term is dominant for $n\gg[t\bar{{\mathcal{J}}}(V_{A_{0}B_{0}})]^{2}$, or $t\ll\sqrt{n}/\bar{{\mathcal{J}}}(V_{A_{0}B_{0}})$.
@@
Finally, we prove
$\bar{{\mathcal{J}}}(V_{A_{0}B_{0}})=MJ$ (90)
to reduce the inequality (V.4.1) to the desired form (76). For this purpose, we generally consider a product state in the form of
@@ V.5 Optimality of the threshold: unbounded entanglement rate for $\alpha<1/2$
$\left|E_{\alpha}(\psi_{t})-E_{\alpha}(\psi_{t=0})\right|\geq\frac{1-2\alpha}{1-\alpha}\log\left(\mathcal{D}_{0}\right)+\frac{1}{1-\alpha}\log\left[\alpha\left(\frac{\bar{{\mathcal{J}}}(V_{A_{0}B_{0}})t}{2}\right)^{2\alpha}\right]$ (94)
under an appropriate initial state $\psi_{t=0})$ [see Eq. (104) below], where we assume
@@
From the above bound, as long as $\alpha<1/2$, the average entanglement rate can be made arbitrarily large regardless of $\bar{{\mathcal{J}}}(V_{A_{0}B_{0}})=1$. Thus, information about the boundary interaction alone does not suffice to upper-bound the Rényi entanglement for $\alpha<1/2$. This point is compared to the geometrically local interacting systems, which will be discussed in Sec. V.5.
@@ V.5.1 Proof of Proposition 8
where we choose as $\varepsilon J_{A}=\varepsilon J_{B}=\pi/2$. Also, from Eq. (97), we get
@@
$=\cos\left(J\Delta t\right)|\psi_{0}\rangle+\sin\left(J\Delta t\right)|1_{A_{1}},2_{A_{0}},2_{B_{0}},1_{B_{1}}\rangle$ (105)
@@
We here note that the condition (99) yields
@@
Here, we consider the following question: “Does a small spectral entanglement rate imply an efficient approximation with low Schmidt rank?” In general, the answer to this question is no. Consequently, we obtain a complexity separation—measured in terms of the Schmidt-rank required for a given approximation error—between approximating the time-evolution unitary and approximating time-evolved states.
@@
Then, one can find a dynamics $e^{-iH_{AB}t}$ such that the error cannot be reduced for any operator $U_{AB,D}$ with a finite Schmidt-rank truncation:
@@ VI.1 Proof of Proposition 9
$\geq t\inf_{W_{2D}}\left\|\tilde{V}_{A_{0}B_{0}}-W_{2D}\right\|-\left(e^{t}-1-t\right)$ (121)
@@
where $W_{s,s^{\prime}}\in\mathbb{R}$ for $\forall s,s^{\prime}$.
@@
The above quantity is known to be equal to the Kolmogorov width of the octahedron [35, 36, 37, 64], denoted by $d_{2D}(B^{N}_{1},\ell_{\infty}^{N})$ (see also Sec. VI.1.1 for a detailed review). For $d_{D}(B^{N}_{1},\ell_{\infty}^{N})$, from Refs. [35] and [37, Proposition 10.10 therein], the Kolmogorov width obeys the inequality of
@@
Note that the lower bound $d_{D}(B^{N}_{1},\ell_{\infty}^{N})=1/2$ is achieved by choosing $W_{s,s^{\prime}}=1/2$ for $\forall s,s^{\prime}$, which has rank $1$.
@@ VI.1.1 A short review on the Kolmogorov width
We here consider $d_{D}(B^{N}_{1},\ell_{\infty}^{N})$ instead of $d_{2D}(B^{N}_{1},\ell_{\infty}^{N})$ for simplicity of notations. We introduce the unit ball $B^{N}_{1}$ as follows:
@@
$=\sum_{j=1}^{D}\vec{w}^{(j)}\vec{w}^{(j)\mathrm{T}}\vec{x}^{\prime}=W\vec{x}^{\prime}$ (139)
@@ VI.2 Conjecture on the operator low-Schmidt-rank approximation
As shown in Proposition 9, the condition that $V_{AB}$ (as well as $e^{-iH_{AB}t}$) has small SE strength (i.e., $\bar{{\mathcal{J}}}(V_{AB})=1$) does not by itself guarantee that the time-evolution operator can be efficiently approximated by a low Schmidt-rank operator. It therefore remains an important open problem to determine under what additional conditions such efficient operator approximations become possible. To address this issue, we introduce an extension of the SE strength, namely the $\alpha$-SE strength $\bar{{\mathcal{J}}}_{\alpha}$ (Definition 2), and conjecture that having small $\alpha$-SE strength provides a sufficient condition for efficient low-rank approximation of the unitary dynamics $e^{-iH_{AB}t}$ (Conjecture 2 below).
@@
However, in many applications, the relevant notion of approximation is the operator norm rather than the Frobenius norm; for instance, in the case of unitary time-evolution operators $e^{-iH_{AB}t}$, the operator norm controls the worst-case state error relevant to complexity and simulation. The approximation properties with respect to the Frobenius norm and the operator norm are qualitatively different. For example, the quantum Fourier transform is known to exhibit maximal operator entanglement when viewed through the lens of the operator Schmidt decomposition [65, 66, 67], yet in terms of the operator norm, there exist efficient low-rank approximations [62]. This discrepancy explains why the operator Schmidt decomposition, while mathematically natural, is insufficient for analyzing low-rank approximations of dynamics in the operator norm.
@@
Finally, without ancillas, the conjecture is trivially false. In this case, the swap operator $S_{AB}$ maps between $A$ and $B$ serves as a counterexample. It creates no entanglement but cannot be approximated by a finite Schmidt rank, as can be shown using similar arguments to (136).
@@ VI.3.1 Stronger assumption
Let us decompose $V_{AB}$ into operator bases $\{V_{j}\}_{j}$, each having small Schmidt rank:
@@
where $D_{0}$ is a constant of order $\mathcal{O}(1)$. Then there exist constants $C_{0}$ and $\kappa$ such that
@@
where $g_{j}\leq\left\|V_{j}\right\|$ and $\{|\phi_{A,j}\rangle,|\phi_{B,j}\rangle\}_{j}$ are normalized states on $A$ and $B$ that are not orthogonal to each other, respectively. Then, let us define the Schmidt decomposition of $V_{AB}|\phi_{0}\rangle$ as
@@
The above inequality is straightforwardly derived from the matrix inequality as [68, 69, 70]
@@
We then set $N_{m}=2^{m}$ and rewrite $\sum_{j}g^{2\alpha}_{j}$ as double sums, which leads to the following inequalities:
@@ VI.3.2 Result on the low-rank approximation
where $\{V_{j}\}_{j}$ are operator bases that constitute the boundary interaction terms as in Eq. (154). Then, under Assumption 10, both for the unitary time evolution $e^{-iHt}$ and the imaginary time evolution $e^{\beta H}$, we can construct approximate operators $\tilde{U}_{t}$ and $\tilde{\rho}_{\beta}$ such that
@@
Remark. The theorem shows that the required Schmidt rank grows only polynomially with respect to the inverse error $1/\epsilon_{0}$, with an exponent proportional to $t{\mathcal{Q}}_{0}$. Importantly, the result guarantees efficient operator approximation provided that the generator $V_{AB}$ itself admits a suitable low-rank decomposition. Although Assumption 10 may appear rather strong at first sight, we will show in Lemma 20 that it is in fact satisfied by a broad class of physically relevant systems, including long-range or power-law decaying interactions. This observation serves as a key backbone for the proof of Theorem 7, where we establish polynomially efficient approximations of long-range quantum Gibbs states.

### v2

@@ III.3 Spectral-entangling strength
A central problem in this work is efficiently characterizing the capacity of an operator to generate entanglement spectra across subsystems. However, within existing methodologies, there has not been an established quantity that efficiently and directly quantifies the ability of a general operator to generate the entanglement spectrum. To address this gap, we introduce a new quantity, which we refer to as the Spectral-Entangling (SE) Strength.
@@
Definition 1 (Spectral-Entangling (SE) Strength).
@@ IV.1 Decay of the Schmidt Coefficients
$\langle A_{j^{\prime}}|A_{j}\rangle\neq 0,\quad\langle B_{j^{\prime}}|B_{j}\rangle\neq 0\quad\textrm{for}\quad j\neq j^{\prime}.$ (21)
@@ IV.1.1 Proof of Lemma 1
Proof of Corollary 3 Using Eq. (29) for each $\lambda_{s}(\phi)$, and applying the definition of the SE strength, we deduce the inequality
@@
Taking the square root on both sides yields the desired bound. $\square$
@@ V.1 Entanglement rate for Rényi entanglement ($\alpha\geq 1/2$)
$E_{\alpha}(\psi_{t})=\frac{1}{1-\alpha}\log\left[{\rm tr}_{A}\left(\rho_{t,A}^{\alpha}\right)\right],\quad\rho_{t,A}={\rm tr}_{B}\left(|\psi_{t}\rangle\langle\psi_{t}|\right),$ (49)
we obtain the upper bound on the entanglement rate as
@@
The statement remains valid even when the system is extended by attaching arbitrary ancillas to $A$ and $B$.
@@
where we use the inequality (18) to obtain $\bar{{\mathcal{J}}}(V_{AB})\leq\sum_{j}|J_{j,AB}|$.
@@ V.2 Optimality of the entanglement generation
Before showing the proof, we emphasize the optimality of the entanglement rate bound for Rényi entanglement. The main points are listed below:
@@
We also make two remarks on optimality:
@@ V.3 Proof of Theorem 1
$\frac{d}{dt}E_{\alpha}(\psi_{t})=\frac{1}{1-\alpha}\frac{d}{dt}\log\left[{\rm tr}_{A}\left(\rho_{t,A}^{\alpha}\right)\right]=\frac{\alpha}{(1-\alpha){\rm tr}\left(\rho_{t,A}^{\alpha}\right)}{\rm tr}_{AB}\left(\rho_{t,A}^{\alpha-1}[-iH,\rho_{t}]\right),$ (59)
where $\rho_{t}=|\psi_{t}\rangle\langle\psi_{t}|$. Precisely speaking, the trace should be restricted to the support of $\rho_{A}$, i.e., the subspace corresponding to strictly positive eigenvalues. In particular, contributions from eigenstates with zero eigenvalues are excluded so that the operator $\rho_{A}^{\alpha-1}$ is well-defined. Then, we obtain
@@
In the following, for an arbitrary quantum state $\rho=|\phi\rangle\langle\phi|$, we generally consider the upper bound of $\left|{\rm tr}_{AB}\left(\rho_{A}^{\alpha-1}[H,\rho]\right)\right|$, where $\rho_{A}={\rm tr}_{B}\left(\rho\right)$.
@@
$\tilde{c}_{\alpha}:=\left|\left(2\alpha-1\right)^{1/(2-2\alpha)}-\left(2\alpha-1\right)^{(2\alpha-1)/(2-2\alpha)}\right|,$ (64)
where we set $\tilde{c}_{1/2}=1$. Note that $\left|x-x^{2\alpha-1}\right|$ attains its maximum when $x=(2\alpha-1)^{1/(2-2\alpha)}$.
@@
From Definition 1 for the SE strength, $V_{AB}|\phi_{s,A},\phi_{s,B}\rangle$ can be rewritten as
@@ V.4 Tightness at $\alpha=1/2$
$\left|\frac{E_{1/2}(\psi_{t})-E_{1/2}(\psi_{0})}{t}\right|\geq 2\bar{{\mathcal{J}}}(V_{A_{0}B_{0}})-\frac{2[\bar{{\mathcal{J}}}(V_{A_{0}B_{0}})]^{2}t}{n}$ (76)
@@
In the inequality (76), only the first term is dominant for $n\gg[\bar{{\mathcal{J}}}(V_{A_{0}B_{0}})]^{2}t$, or $t\ll n/[\bar{{\mathcal{J}}}(V_{A_{0}B_{0}})]^{2}$.
@@
Finally, we prove the equation
$\bar{{\mathcal{J}}}(V_{A_{0}B_{0}})=MJ,$ (90)
which reduces the inequality (V.4.1) to the desired form (76). For this purpose, we generally consider a product state in the form of
@@ V.5 Optimality of the threshold: unbounded entanglement rate for $\alpha<1/2$
$\left|E_{\alpha}(\psi_{t})-E_{\alpha}(\psi_{t=0})\right|\geq\frac{1-2\alpha}{1-\alpha}\log\left(\mathcal{D}_{0}\right)+\frac{1}{1-\alpha}\log\left[\alpha\left(\frac{\bar{{\mathcal{J}}}(V_{A_{0}B_{0}})t}{2}\right)^{2\alpha}\right],$ (94)
under an appropriate initial state $|\psi_{t=0}\rangle$ [see Eq. (104) below], where we assume
@@
From the above bound, as long as $\alpha<1/2$, the average entanglement rate can be made arbitrarily large regardless of $\bar{{\mathcal{J}}}(V_{A_{0}B_{0}})=1$. Thus, information about the boundary interaction alone does not suffice to upper-bound the Rényi entanglement for $\alpha<1/2$. This point is compared to the geometrically local interacting systems, which will be discussed in Sec. IX.
@@ V.5.1 Proof of Proposition 8
where we choose $\varepsilon J_{A}=\varepsilon J_{B}=\pi/2$. Also, from Eq. (97), we get
@@
$=\cos\left(J\Delta t\right)|\psi_{0}\rangle+\sin\left(J\Delta t\right)|1_{A_{1}},2_{A_{0}},2_{B_{0}},1_{B_{1}}\rangle.$ (105)
@@
Here, we note that the condition (99) yields
@@
Here, we consider the following question: “Does a small spectral entanglement rate imply an efficient approximation with low Schmidt rank?” In general, the answer to this question is no. Consequently, we obtain a complexity separation—measured in terms of the Schmidt‑rank required for a given approximation error—between approximating the time‑evolution unitary and approximating time‑evolved states.
@@
Then, there always exists a dynamics $e^{-iH_{AB}t}$ such that the error cannot be reduced for any operator $U_{AB,D}$ with a finite Schmidt-rank truncation:
@@ VI.1 Proof of Proposition 9
$\geq t\inf_{W_{2D}}\left\|\tilde{V}_{A_{0}B_{0}}-W_{2D}\right\|-\left(e^{t}-1-t\right).$ (121)
@@
where $W_{s,s^{\prime}}\in\mathbb{R}$ for all $s,s^{\prime}$.
@@
The above quantity is known to be equal to the Kolmogorov width of the octahedron [35, 36, 37, 64], denoted by $d_{2D}(B^{N}_{1},\ell_{\infty}^{N})$ (see also Sec. VI.1.1 for a detailed review). For $d_{D}(B^{N}_{1},\ell_{\infty}^{N})$, from Refs. [35] and [37, Proposition 10.10 therein], the Kolmogorov width obeys the inequality
@@
Note that the lower bound $d_{D}(B^{N}_{1},\ell_{\infty}^{N})=1/2$ is achieved by choosing $W_{s,s^{\prime}}=1/2$ for all $s,s^{\prime}$, which has rank $1$.
@@ VI.1.1 A short review on the Kolmogorov width
We now consider $d_{D}(B^{N}_{1},\ell_{\infty}^{N})$ instead of $d_{2D}(B^{N}_{1},\ell_{\infty}^{N})$ for simplicity of notation. We introduce the unit ball $B^{N}_{1}$ as follows:
@@
$=\sum_{j=1}^{D}\vec{w}^{(j)}\vec{w}^{(j)\mathrm{T}}\vec{x}^{\prime}=W\vec{x}^{\prime},$ (139)
@@ VI.2 Conjecture on the operator low-Schmidt-rank approximation
As shown in Proposition 9, the condition that $V_{AB}$ (as well as $e^{-iH_{AB}t}$) has small SE strength (i.e., $\bar{{\mathcal{J}}}(V_{AB})=1$) does not by itself guarantee that the time‑evolution operator can be efficiently approximated by a low Schmidt-rank operator. It therefore remains an important open problem to determine under what additional conditions such efficient operator approximations become possible. To address this issue, we introduce an extension of the SE strength, namely the $\alpha$-SE strength $\bar{{\mathcal{J}}}_{\alpha}$ (Definition 2), and conjecture that having small $\alpha$-SE strength provides a sufficient condition for efficient low-rank approximation of the unitary dynamics $e^{-iH_{AB}t}$ (Conjecture 2 below).
@@
However, in many applications, the relevant notion of approximation is the operator norm rather than the Frobenius norm; for instance, in the case of unitary time-evolution operators $e^{-iH_{AB}t}$, the operator norm controls the worst-case state error relevant to complexity and simulation. The approximation properties with respect to the Frobenius norm and the operator norm are qualitatively different. For example, the quantum Fourier transform is known to exhibit maximal operator entanglement when viewed through the lens of the operator Schmidt decomposition [65, 66, 67]; with respect to the operator norm, however, efficient low‑rank approximations exist [62]. This discrepancy explains why the operator Schmidt decomposition, while mathematically natural, is insufficient for analyzing low-rank approximations of dynamics with respect to the operator norm.
@@
Finally, in the absence of ancillas, the conjecture is trivially false. In this case, the swap operator $S_{AB}$ maps between $A$ and $B$ serves as a counterexample. It creates no entanglement but cannot be approximated by a finite Schmidt rank, as can be shown using similar arguments to (136).
@@ VI.3.1 Stronger assumption
Given the decomposition of $V_{AB}$ into operator bases $\{V_{j}\}_{j}$, each having small Schmidt rank:
@@
where $D_{0}$ is a constant of order $\mathcal{O}(1)$. Then, there exist constants $C_{0}$ and $\kappa$ such that the following inequality holds:
@@
where $g_{j}\leq\left\|V_{j}\right\|$ and $\{|\phi_{A,j}\rangle,|\phi_{B,j}\rangle\}_{j}$ are normalized states on $A$ and $B$, respectively, that are not orthogonal to each other. Next, we consider the Schmidt decomposition of $V_{AB}|\phi_{0}\rangle$ as
@@
The above inequality follows directly from the matrix inequality as [68, 69, 70]
@@
We set $N_{m}=2^{m}$ and rewrite $\sum_{j}g^{2\alpha}_{j}$ as double sums, which leads to the following inequalities:
@@ VI.3.2 Result on the low-rank approximation
where $\{V_{j}\}_{j}$ are operator bases that constitute the boundary interaction terms as in Eq. (154). Then, under Assumption 10, for both the unitary time evolution $e^{-iHt}$ and the imaginary time evolution $e^{\beta H}$, we can construct approximate operators $\tilde{U}_{t}$ and $\tilde{\rho}_{\beta}$ such that
@@
Remark. The theorem shows that the required Schmidt rank grows only polynomially with respect to the inverse error $1/\epsilon_{0}$, with an exponent proportional to $t{\mathcal{Q}}_{0}$. Importantly, the result guarantees efficient operator approximation provided that the generator $V_{AB}$ itself admits a suitable low-rank decomposition. Although Assumption 10 may appear rather strong at first sight, we will show in Lemma 20 that it is in fact satisfied by a broad class of physically relevant systems, including long-range or power-law decaying interactions. This observation serves as a key ingredient for the proof of Theorem 7, where we establish polynomially efficient approximations of long-range quantum Gibbs states.